(1 point) a. Find the most general real-valued solution to the linear system of differential equations * = [- xi(t) = C1 + CI AM REALLY STRUGGLING ON THIS PROBLEM PLEASE HELP ME CORRECT AND NEAT WORK IS MUCH APPRECIATED THANKS
Answer #1

a) XE=; -G o 2 3 -5 Y- २ In-del=0 => -3-5-1 => 12+54+6=0 d=-2 -2=-3 d=-2 [A+QI]x = 0 2 2 -3 -3 J[3]= x=-4 v Fi] d=-3 -O [A+31]x = 3 a -2 2 1-13) = -3 A= -24 3 let y=3 = -2 V = 3 2 X(E-) = Get [11]+sé - Pocit) [SCE), 13] = 4 -QE7 -at + -2b) here eigenvalues -2,-3 are reals and both are negative and distinct so in phase plane system is sink/stable node

Ans) sink/stable node

For part a we first find eigenvalues and eigenvectors of Matrix A.